Carbon tetrachloride reacts at high temperatures with oxygen to produce two toxic gases, phosgene and chlorine.
CCl4(g) + 1/2O2(g) COCl2(g) + Cl2(g), Kc = 4.4 × 109 at 1,000 K
Calculate Kc for the reaction 2CCl4(g) + O2(g) 2COCl2(g) + 2Cl2(g).
Answer #1
2/4.4*10^9
=4.5 * 10^-10
Answer #2
CCl4(g) + (1/2)O2(g) <----------> COCl2(g) + Cl2(g)
Kc1 = [Cl2][COCL2]/[CCL4][O2]1/2 = 4.4 x 109

2CCl4(g) + O2(g) <-------> 2COCl2(g) + 2Cl2(g)
Kc2 = [Cl2]2[COCL2]2/[CCL4]2[O2]
= ([Cl2][COCL2]/[CCL4][O2]1/2)2
= (Kc1)2 = (4.4 x 109)2 = 1.9 x 1019

The answer is D. 1.9 x 10^19
answered by: Food4thought...
Answer #3
Kc(1) = P(Cl2).P(COCl2)/[P(CCl4).P(O2)^1/2)
Kc(2) = [P(Cl2)^2.P(COCl2)^2]/[P(CCl4)^2.P(O2)] = Kc(1)^2
= (4.4)^2 x 10^18 = 1.94 x 10^19.

Which is of course sensible as delta.G of 2 is twice delta.G of 1 so Kc is squared.
answered by: jame smith